For a step index fibre with a core of refractive index of 1.54 and cladding refractive index 1.5 calculate the following when fibre is kept in water

a)Fractional refractive index



1
Expert's answer
2021-05-10T12:17:31-0400

Fractional Refractive index, "\Delta=\dfrac{\mu_1-\mu_2}{\mu_1}"

"\Delta=\dfrac{1.54-1.50}{1.54}=0.025"